The Connecticut Computer Company has the selected financial results shown below.

20% Debt 40% Debt 75% Debt
Debt $30000
Equity 120000
Total Capital $150000
Shares@ $5 24000
EBIT $24000
Interest (15%) 4500
EBT $19500
Tax (40%) 7800
Net Income $11700
ROE
EPS

The company is considering a capital restructuring to increase leverage from its present level of 20% of capital.

  1. Calculate Connecticut's ROE and EPS under its current capital structure.
  2. Restate the financial statement line items shown, the number of shares outstanding, ROE, and EPS if Connecticut borrows money and uses it to retire stock until its capital structure is 40% debt assuming EBIT remains unchanged and the stock continues to sell at its book value. (Develop the second column of the chart shown.)
  3. Recalculate the same figures assuming Connecticut continues to restructure until its capital structure is 75% debt. (Develop the third column of the chart.)
    Round ROE to one decimal place. Round EPS to two decimal places. Enter all amounts as positive numbers.
    20% Debt 40% Debt 75% Debt
    Debt $30000 $     $    
    Equity 120000          
    Total Capital $150000 $     $    
    Shares@ $5 24000          
    EBIT $24000          
    Interest (15%) 4500          
    EBT $19500 $     $    
    Tax (40%) 7800          
    Net Income $11700 $     $    
    ROE   %   %   %
    EPS $      $     $    
  4. How is increasing leverage affecting financial performance? What overall effect might the changes have on the market price of Connecticut’s stock? Why? (Words only. Hint: Consider the move from 20% to 40% and that from 40% to 75% separately.)
    The input in the box below will not be graded, but may be reviewed and considered by your instructor.
Answer #1

a.

Calculation of ROE & EPS

no of shares outstanding = $120000/$5 =24000 shares

return on equity = net income available to equity shareholder/ equity capital

= $11700/$120000*100

=9.75%

EPS = net income available to equity shareholder/no of equity shares

= $11700/24000

= $ 0.4875 per share

b. When capital structure is 40% debt

Debt = 40% of $150000 =$60000

Equity = 60% of 150000 = 90000

no of shares = 90000/5 =18000 shares

Note 1. Calculation of net income available to equity shareholder

EBIT 24000
Less: interest(60000*15%) 9000
EBT 15000
Less: Tax @ 40% 6000
Net income 9000

Calculation of ROE & EPS

return on equity = net income available to equity shareholder/ equity capital

= $9000/$90000*100

=10%

EPS = net income available to equity shareholder/no of equity shares

= $9000/18000

= $ 0.5 per share

c. When capital structure is 75% debt

Debt = 75% of $150000 =$112500

Equity = 25% of 150000 = $37500

no of shares = 37500/5 =7500 shares

Note 1. Calculation of net income available to equity shareholder

EBIT 24000
Less: interest(112500*15%) 16875
EBT 7125
Less: Tax @ 40% 2850
Net income 4275

Calculation of ROE & EPS

return on equity = net income available to equity shareholder/ equity capital

= $4275/$37500*100

=11.4%

EPS = net income available to equity shareholder/no of equity shares

= $4275/7500

= $ 0.57 per share

d. Increasing leverage affecting financial performance of co much. when debt funding increase ROE and EPS also increasing proportionately. but some time it may negatively affect the financial performance of company.

Overall effect might change the market price of stock , it may go up as there is increase in EPS & ROE or might decrease due to high risky company.

on an appropriate proportion it may constant and benefit the company on reducing cost of capital employed.